A man can row at 5km/h in stationary water. It takes him 4 hours to row upstream towards a bridge and back to his house. How far is the bridge from the house if the river flows at 2km/h?
Answer by jorel1380(3719)

x/(5+2)+x/(5-2)=4
x/7+x/3=4
3x+7x=84
10x=84
x=8.4kms..